OpenStudy (anonymous):

Which equations are equivalent to the given equation? 0.4y=10 Choose all answers that are correct. A. 0.4y=10÷0.4 B. 0.4y÷0.4=10÷0.4 C. 0.4y*2.5=10*2.5 D. 0.4y÷0.4=10*0.4

OpenStudy (anonymous):

First, what can you do to get y by itself?

OpenStudy (anonymous):

I'm completely lost here.

OpenStudy (anonymous):

Haha np. We're basically trying to find different ways to simplify. If we have 0.4y=10 we can get y by itself by dividing 0.4 Ex: 0.4y/0.4 gets y alone. But what you do to one side, you must do to the other. So 0.4y/0.4 = 10/0.4

OpenStudy (anonymous):

Does that make sense at all?

OpenStudy (anonymous):

So B would be one of the answers?

OpenStudy (anonymous):

That would be correct! So when we solve the equation 0.4y=10 we end up with y=10/0.4 y=25. Do you see any other equations that get the same answer?

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

Well lets try it out. 0.4y = 10/0.4 This simplifies out to be 0.4y = 25 Then, divide 0.4 from y, and (what you do to one side you do to the other) So you're gonna end up with y= 24/0.4 which gives you y=62.5

OpenStudy (anonymous):

So A is incorrect

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

It looks similar, so its kind of tricky

OpenStudy (anonymous):

C is another correct answer. 10*2.5=25

OpenStudy (anonymous):

Good job!! 0.4y x 2.5 = 10 x 2.5 0.4y/2.5 = 10 x2.5/2.5 (You do this to cancel it out) So you're gonna end up with 0.4y = 10 which brings you back to your original equation :)

OpenStudy (anonymous):

I believe A & C are the only correct answers. Am I correct?

OpenStudy (anonymous):

Well remember we proved A to be wrong, because it actually equals 62.5 B and C are the correct ones haha

OpenStudy (anonymous):

Sorry I meant B & C

OpenStudy (anonymous):

If my question is answered and the only correct answers are B & C then I thank you for helping me with my question and I will reward you with a medal.

OpenStudy (anonymous):

I believe those are the only correct ones! We proved A untrue, and D just seems to be all sorts of wrong. So no problem!

OpenStudy (anonymous):

Thank you very much.